OpenStudy (anonymous):

Can I get some help with solving an implicit differentiation problem? : x^4(x+y)=y^2(3x-y)

OpenStudy (dumbcow):

use product rule and remember to multiply by dy/dx every time you differentiate a "y" term \[4x^3 (x+y) + x^4(1 + \frac{dy}{dx}) = 2y \frac{dy}{dx}(3x-y) + y^2 (3-\frac{dy}{dx})\]

OpenStudy (anonymous):

alright, thank you for your help

OpenStudy (dumbcow):

can you solve for dy/dx from here

OpenStudy (anonymous):

yes, I can. though can you show me how you used the product rule to turn \[x^4(x+y)\] into \[4x^3(x+y)+x^4(1+\frac{ dy }{ dx }\]

OpenStudy (anonymous):

because I got everything except that (1+...) in the second term

OpenStudy (dumbcow):

derivative of (x+y) = 1 + dy/dx

OpenStudy (anonymous):

oh, haha thanks xD I forgot about the x

OpenStudy (solomonzelman):

yes, dumbcow, tnx for doing user's problem for him/her -:(

OpenStudy (anonymous):

Refer to the attachment from Mathematica 9.